# Vision-Language Models as a Source of Rewards

### 14 Dec 2023 v1

@@ 2.1 VLMs Rewards
A number of prior works have investigated using CLIP as a reward function for RL. Most closely related to our work is MineDojo (Fan et al., 2022) which first finetunes CLIP with Minecraft videos to form MineCLIP. Then MineCLIP is used as a dense shaping reward function in addition to a ground truth binary reward function for programmatic tasks where ground truth is available, except for creative tasks where no ground truth is available. The main differences between this work and MineDojo are that (i) we use an off-the-shelf CLIP model, (ii) we do not utilize any ground truth information during RL training. and (iii) we train a single RL agent that can solve thousands of language-based goals. Similar to MineDojo, CLIP4MC (Ding et al., 2023) finetunes CLIP on a curated dataset of Minecraft Youtube videos. CLIP4MC then uses PPO as the RL training backbone and the policy is trained on the CLIP4MC reward in the same form as MineCLIP reward together with the sparse completion where available. CLIP has also been used for reward-shaping simulated robotics domain (Cui et al., 2022; Mahmoudieh et al., 2022; Sontakke et al., 2023). RoboCLIP (Sontakke et al., 2023) computes a sparse trajectory level reward as the cosine similarity score between the embedding of the trajectory video and the embedding of the text description of the task.
CLIP encoder was used for semantic intrinsic reward signal for curiosity-driven exploration in (Tam et al., 2022). An orthogonal approach for scaling up reward modeling is to use Large Language Models (LLMs) to translate task specification into reward functions executable programs, resulting in sparse rewards (Yu et al., 2023) or dense rewards (Xie et al., 2023). Earlier works also use LLMs itself as binary reward in negotiation games (Kwon et al., 2023), collaborative human-AI interaction games (Hu and Sadigh, 2023), or dense reward based on the cosine similarity of a caption (hard-coded or learned) of transition and goals in a 2D Crafter environment (Du et al., 2023b; Hafner, 2021).
@@ 6 Authors and Contributions
We list authors alphabetically by last name. Please direct all correspondence to Volodymyr Mnih (vmnih@google.com), Feryal Behbahani (feryal@google.com), Michael Laskin (mlaskin@google.com) and Harris Chan (harrischan@google.com).

### 22 Feb 2024 v2

@@ 2.1 VLMs Rewards
A number of prior works have investigated using pretrained models such as CLIP as a reward function for RL. Most closely related to our work is MineDojo (Fan et al., 2022) which first finetunes CLIP with Minecraft videos to form MineCLIP. Then MineCLIP is used as a dense shaping reward function in addition to a ground truth binary reward function for programmatic tasks where ground truth is available, except for creative tasks where no ground truth is available. The main differences between this work and MineDojo are that (i) we use an off-the-shelf CLIP model, (ii) we do not utilize any ground truth information during RL training. and (iii) we train a single RL agent that can solve thousands of language-based goals. Similar to MineDojo, CLIP4MC (Ding et al., 2023) finetunes CLIP on a curated dataset of Minecraft Youtube videos. CLIP4MC then uses PPO as the RL training backbone and the policy is trained on the CLIP4MC reward in the same form as MineCLIP reward together with the sparse completion where available. CLIP has also been used for reward-shaping simulated robotics domain (Cui et al., 2022; Mahmoudieh et al., 2022; Sontakke et al., 2023). RoboCLIP (Sontakke et al., 2023) computes a sparse trajectory level reward as the cosine similarity score between the embedding of the trajectory video and the embedding of the text description of the task. VLM-RMs (Rocamonde et al., 2023) proposes goal-baseline regularization by projecting out irrelevant CLIP embedding space dimensions from the goal embedding using a secondary baseline text prompt embedding, applied to training a controller for a simulated humanoid robot. CLIP encoder was used for semantic intrinsic reward signal for curiosity-driven exploration in (Tam et al., 2022). VIPER (Escontrela et al., 2023) uses the conditional log-likelihood of a frozen action-free video prediction model, such as VideoGPT (Yan et al., 2021), summed with an entropy bonus as the reward signal for reinforcement learning agents without groundtruth task rewards in DeepMind Control Suite (Tunyasuvunakool et al., 2020), Atari (Bellemare et al., 2012), and RLBench (James et al., 2020) tasks.
An orthogonal approach for scaling up reward modeling is to use Large Language Models (LLMs) to translate task specification into reward functions executable programs, resulting in sparse rewards (Yu et al., 2023) or dense rewards (Xie et al., 2023; Ma et al., 2023). Earlier works also use LLMs itself as binary reward in negotiation games (Kwon et al., 2023), collaborative human-AI interaction games (Hu and Sadigh, 2023), or dense reward based on the cosine similarity of a caption (hard-coded or learned) of transition and goals in a 2D Crafter environment (Du et al., 2023b; Hafner, 2021).
@@ 6 Authors and Contributions
We list authors alphabetically by last name. Please direct all correspondence to Volodymyr Mnih (vmnih@google.com), Feryal Behbahani (feryal@google.com), and Harris Chan
(harrischan@google.com).